hero (hero):

Rubik's cube group is an algebraic group with elements corresponding to the cube operations on the toy called the Rubik's cube. Discuss the formal description of how Rubik's cube could correspond to an algebraic group and use the methods of group theory to determine the order of the group (or in this context, the number of ways to arrange the color squares on the Rubik's cube).

OpenStudy (kinggeorge):

At my current level of knowledge, I'd probably try to somehow find the orbits and conjugacy classes of the Rubik's Cube. If we can find those, and the orders of these, then it's rather easy to find the total number of arrangements possible.

OpenStudy (kinggeorge):

The hardest part of that idea, would probably be finding the conjugacy classes.

hero (hero):

Would you like to discuss this with me in vyew? I have compiled some information that might help the both of us figure this out. I already know the total number of arrangements possible. I just need to know how they got it. I have plenty of information on it and I can piece it all together for you and maybe you can help me come up with a competent, logical explanation the links all of these elements together

OpenStudy (kinggeorge):

I'm really sorry, but I actually have to leave now, but when I get back in about 3-4 hours I would be happy to.

hero (hero):

Okay, looking forward to it.

hero (hero):

I've already found a resource with the conjugacy classes already calcuated, so now, I just need help with using that to explain how to find the order of the group

hero (hero):

@myininaya @satellite73 @Zarkon @JamesJ @dumbcow @FoolForMath @asnaseer

hero (hero):

@amistre64

hero (hero):

@anyone

OpenStudy (kinggeorge):

It took a little bit longer than expected, but I'm back now if you want to discuss this now. BTW, where is that resource with the conjugacy classes enumerated?

hero (hero):

Yeah, I have it but I think there's something wrong with it. Would you like to go to vyew to discuss this?

OpenStudy (kinggeorge):

Sure. I need to sign up for that right?

hero (hero):

No, I will post the link

hero (hero):

I sent you the link

hero (hero):

Hey @KingGeorge I found the section in that pdf document that talks about using cube moves as group elements. This is perfect!

OpenStudy (kinggeorge):

I'm glad it was so helpful :P

hero (hero):

I still don't have a good solution yet.

OpenStudy (chaise):

I'd like to look at your work if you don't mind. This seems interesting.

hero (hero):

I don't have any useful work at the moment.

hero (hero):

If I did, I wouldn't be up at 4am

OpenStudy (chaise):

Anything and everything that might keep my mind occupied?

hero (hero):

Oh, I have plenty of related links to share if you think you can piece something together.

OpenStudy (chaise):

Sure - post em, I don't like your chances though. Just heavily intrigued by the whole idea.

hero (hero):

I don't know why you're intrigued by it. Honestly, this assignment disgusts me, but if you're as interested as you say you are, you have a better chance than me of piecing something together.

hero (hero):

I will post some links soon

OpenStudy (chaise):

I'd even be curious to see your task sheet, if that is what is handed out for the assignment. What questions are postulated and posed, etc, etc,

hero (hero):

The question above is the only task. I don't know what you mean by a "task sheet"

OpenStudy (chaise):

Ah okay - for assessments I'm given a sheet with the criteria for marking, etc, etc.

hero (hero):

There's no restrictions on a response to this by the way.

hero (hero):

No format restrictions.

hero (hero):

I will post some links now.

hero (hero):

I will post one of the responses made by another person even though it isn't completely correct because she didn't show directly how the order can be derived from the operations of the Rubik Group.

hero (hero):

Sample Response (although not completely correct): In order for an algebraic structure to be a group, it must be closed under the operation, it must contain an identity element, every element must have an inverse, and it must be associative. By this definition, the possible moves of the Rubik’s cube can be represented as a group. Let (G, *) be the group of possible moves. If M1 and M2 are both moves that can be made on the cube, then M1 * M2, which essentially means complete move one then complete move two, must also be a move. (G, *) does have an identity element called e which, in the case of the Rubik’s cube, represents an empty move. M * e is essentially the same things as completing move one and then doing nothing, so M * e = M. Logically, every possible move in (G, *) must have an inverse, or a move that can undo the original move. Completing the move and then undoing the move is the same thing as doing nothing, so M * M’ = e. To show that (G, *) is associative, First, let’s investigate what a sequence of two moves does to the cubie. If M1 and M2 are two moves, then M1 * M2 is the move where we first do M1 and then do M2. The move M1 moves C to the cubicle M1(C); the move M2 then moves it to M2(M1(C)). Therefore, (M1 * M2)(C) = M2(M1(C)). To show that * is associative, we need to show that (M1 * M2) * M3 = M1 * (M2 * M3) for any moves M1, M2, and M3. This is the same as showing that (M1 * M2) * M3 and M1 * (M2 * M3) do the same thing to every cubie. That is, we want to show that [(M1 * M2) * M3](C) = [M1 * (M2 * M3)](C) for any cubie C. We know from our above calculation that [(M1 * M2) * M3](C) = M3([M1 * M2](C)) = M3(M2(M1(C))). On the other hand, [M1 * (M2 * M3)](C) = (M2 * M3)(M1(C)) = M3(M2(M1(C))). So, (M1 * M2) * M3 = M1 * (M2 * M3). Thus, * is associative. (Chen, n.d., p. 11) Since (G, *) fits all the properties, it is a group. Although the order of this group is large, it is still finite. There are 2^(12)3^(8)8!12! possible ways to configure the Rubik’s cube. Not all of these configurations are considered valid because they cannot be achieved by a series of moves from a starting configuration (Chen, n.d.). There are 4 x 10^19 valid possible ways to configure the Rubik’s cube (Chen, n.d.). References Chen, J. (n.d.). Group theory and the Rubik’s cube. Retrieved from http://www.math.harvard.edu/~jjchen/docs/Group%20Theory%20and%20the%20Rubik's%20Cube.pdf

OpenStudy (chaise):

Correct me if I'm wrong but the only way you can prove associative is through isomorphism? A bit of an amateur here. I may be wrong.

hero (hero):

I wouldn't assume that isomorphism is the only way to prove associative.

hero (hero):

If anyone else wants to help me out additionally, feel free.

OpenStudy (anonymous):

this is easy

OpenStudy (anonymous):

i could solve a rubik's cube blindfolded :P

OpenStudy (turingtest):

@Tai so can I, but that's got absolutely nothing to do with the question. please keep your comments relevant.